Find the indicated probabilities using the geometric​ distribution, the Poisson​ distribution, or the binomial distribution. Then determine if the events are unusual. If​ convenient, use the appropriate probability table or technology to find the probabilities.

Fifty dash sevenFifty-seven

percent of parents would give up cable television to have their​ child's education paid for. You randomly select

fivefive

parents. Find the probability that the number of parents who would give up cable television to have their​ child's education paid for is​ (a) exactly

twotwo​,

​(b) less than

fourfour​,

and​ (c) at least

threethree.
Answer #1
here this is binomial with parameter n=5 and p=0.57

a)

P(X=2)= (nCx)px(1−p)(n-x)    = 0.2583 (please try 0.258 if required to 3 decimals)

b)

P(X<4)= x=0a     (nCx)px(1−p)(n-x)    = 0.7129 (please try 0.713 if required to 3 decimals)

c)

P(X>=3)=1-P(X<=2)= 1-∑x=0x-1   (nCx)px(q)(n-x) = 0.6295 (please try 0.630 if required to 3 decimals)